OWG Meeting 26-Mar-2014

From GlueXWiki
Jump to: navigation, search

Location and Time

Room: CC F326-327

Time: 1:30pm-2:30pm

Remote Connection

ESNet: 8542553

(if problems, call phone in conference room: 757-269-6460)


Bluejeans Information

The meeting this week will be made using the bluejeans video conferencing solution supported by Jefferson Lab. You can find information from Jefferson Lab on this product as well as some notes that we within GlueX have made GlueX Bluejeans Video Conferencing. You should be able to connect seamlessly with virtually any device. If you have a polycom device, or want to connect through your web browser or tablet, use the Network Connection below. If you want to connect via phone, then use the phone information from below. The main drawback is that you may not be able to connect via a browser using a linux computer.

Bluejeans Applications

The first time that you connect to a bluejeans meeting, you will be prompted to install an application for your web browser. Versions of these for windows and macos have been placed on docDb as document 2435. You can also find apps for iOS and android.

Connect to the Meeting

Join via Polycom room system

    • Network Connection
      • IP Address:
      • URL: bjn.vc
    • Enter Meeting ID: 462509408

Join via a Webbrowser

    • If you have not registered with bluejeans through JLab:
    • If you have registered with bluejeans through JLab:
      • URL: [2]
      • Enter Conference ID: 462509408

Join via Phone

    • Dial-in phone number:
      • US or Canada: +1 408 740 7256 or
      • US or Canada: +1 888 240 2560
      • Dial-in international phone numbers
    • Enter Conference ID: 462509408

Previous Meeting


  • Announcements
  • Major systems reports
    • Trigger
    • DAQ
    • Monitoring
    • Controls
  • Gluon filesystem structure and code versioning (Sergey F.)
  • Mantis Task Tracker


Attendees: David L.(chair), Simon T., Chris C., Beni Z., Sergey P., Hovanes E., Curtis M., Sean D., Graham H., Dave A., Mark D., Dmitry R., Vardan G., Carl T., Nerses, Alex S., Sergey F., Bryan M., Eugene C.



  • Last week started modifying step to include TS, but keep the TI-based DAQs.
  • Required adding a second optical transceiver to the TI boards and switching the one used by stand-alone DAQ so that TS could use the one in port 1.
  • Lots of different issues. These seem to be related to adding the second optical transceiver.
  • All systems can run with internal pulser (TI-based)
  • Have not been able to check out TS-based system with multiple transceivers
  • Port 1 seems to work, but Port 5 does not
    • May need to temporarily switch TI back to port 1 in order to get it working.
  • Chris C. reports that cabling for the FDC is complete.
  • CDC and FC will not have stand-alone DAQ systems but will only run in TS mode.
    • This is plan for now, but TI-base DAQ can be set up if needed later


  • DAQ group has implemented several bug fixes for CODA. These are now installed in CODA 3.01 directory
  • Work continues on finalizing format of configuration files for trigger and and DAQ
    • This is nearly doe, but has been delayed slightly by higher priority DAQ/Trigger issues


  • Issue with the JANA EVIO parser that caused 10 of the 12 modules read out by the BCAL to be missed is resolved
    • Problem was in format of filler word from f250ADC that changed in 2013.
  • Farm controller development continues. Dave and Vardan have been communicating regularly
  • There was some discussion of the rocid values to be used for production running. The Translation Table in the CCDB uses on convention, but the trigger group has started using a different convention.
    • we will come to some consensus offline and implement it (John L., Alex S., Mark D., and David L.)


  • Dave B. is working on BCAL power supply and chiller interlocks
  • Orlando:
    • Working on BCAL pulser control GUI (nearly finished)
    • Bios for the LEDs is still left
    • BCAL group is not using quite yet
  • A temporary solution for backing up and restoring CDC and FDC voltages has been implemented
  • Archiving is not yet active


  • Sergey F. has been looking at organization of the gluon system. In particular, a directory structure that can be stored in a version control system, and checked out into individual user accounts and used
  • Sergey suggested we consider whether it makes sense to have a repository on the gluon cluster. The alternative is to use the Hall-D repository where the rest of the Hall-D specific code resides. People were asked to consider this so it could be discussed next time.
  • Alex suggested we consider creating separate user accounts for different detector systems that could help maintain a separation of files used for configuration etc. from the general hdops account (which is quite fluid). We agreed to consider it and discuss it at the next meeting in the context of Sergey's presentation.

Next Meeting

Next meeting is scheduled for April 9th